OpenStudy (anonymous):

At a certain time of day, a tree that is x meters tall casts a shadow that is x − 46 meters long. If the distance from the top of the tree to the end of the shadow is x + 4 meters long, what is the height, x, of the tree?

OpenStudy (misty1212):



OpenStudy (misty1212):

hi!!

OpenStudy (misty1212):

i guess you are supposed to use pythgagoras and solve \[x^2+(x-46)^2=(x+4)^2\]

OpenStudy (anonymous):

Right that is correct. I have 70 and 30 for answers. I don't know if that is right or which one is right

OpenStudy (misty1212):

you need \(x\) oh i see lets check

OpenStudy (anonymous):

I think 30 is the right answer?

OpenStudy (misty1212):

i think they are both right except for 30 which is wrong

OpenStudy (misty1212):

i mean they both are solutions to the original equation, but it is unlikely that the shadow is \(30-46=-16\) feet long isn't it?

OpenStudy (anonymous):

Yes, 30 is wrong as the shadow is x-46m long. Which doesn't make sense :P

OpenStudy (anonymous):

HAHA! True! I guess I should probably use my head a little more! Thanks so much!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]